OpenStudy (anonymous):

36x^2+85x+49=x solve by factoring

OpenStudy (anonymous):

36x^2 + 85x + 49 = x Minus x from both sides of the equation 36x^2 + 84x + 49 = 0 Then multiply 36 by 49 and you get 1764 Find two numbers that multiply to be 1764 but add up to be 84

OpenStudy (campbell_st):

well the 1st and last terms are perfect squares and if you subtract x from both sides the coefficient of the middle term changes to 84 which is double the product of the terms in a perfect square. hope this helps.. as you are solving \[36x^2 + 84x + 49 = 0\]

OpenStudy (campbell_st):

ps... don't worry about multiplying anything... what squared is 49, what squared isn 36x^2

OpenStudy (anonymous):

im so lost

OpenStudy (campbell_st):

ok... this is a perfect square \[(a + b)^2 = a^2 + 2ab + b^2\] your question is \[36x^2 + 84x + 49 = 0\] can you find the values of a and b... in your equation...?

OpenStudy (anonymous):

idk honestly

OpenStudy (campbell_st):

ok... what is \[\sqrt{36} =......... \sqrt{x^2}=..............\sqrt{49}=.............\]

OpenStudy (campbell_st):

any ideas yet..?

OpenStudy (anonymous):

Hey @higherlearning r u even there??

OpenStudy (anonymous):

Luk just follow wat u did in d previous sum.. Remember d way v did it? @higherlearning ??

OpenStudy (anonymous):

okay let me help you out.. okay?? @higherlearning ?

OpenStudy (anonymous):

36x^2+85x+49=x now subtact x from both sides 36x^2+85x+49-x=x-x so we get... 36x^2+84x+49=0 now we need to find 2 such factors of 36 X 44 which must add upto or must subtracted to obtain 1764 which is (36 X 44)... now the factors are 42 and 42.. can u solve it now?

OpenStudy (anonymous):

in case u dont and reply"i'm so lost"! lol! here's how.. \[36x^2+42x+42x+49=0\] \[6x(6x+7x)+7x(6x+7x)=0\] \[(6x+7)(6x+7)=0\] so its \[6x+7=0 \] or \[6x=-7\] or \[x=-7\div 6\] okay??